# An Overview of Cyber Security Funding for Open Source Software

### v1

@@ Abstract
Many open source software (OSS) projects need more human resources for maintenance, improvements, and sometimes even their survival. This need allegedly applies even to vital OSS projects that can be seen as being a part of the world’s critical infrastructures. To address this resourcing problem, new funding instruments for OSS projects have been established in recent years. The paper examines two such funding bodies for OSS and the projects they have funded. The focus of both funding bodies is on software security and cyber security in general. Based on a qualitative analysis, particularly OSS supply chains, network and cryptography libraries, programming languages, and operating systems and their low-level components have been funded and thus seen as critical in terms of cyber security by the two funding bodies. In addition to this and other results, the paper makes a contribution by connecting the research branches of critical infrastructure and sustainability of OSS projects. A further contribution is made by connecting the topic examined to recent cyber security regulations. Furthermore, an important argument is raised that neither cyber security nor sustainability alone can entirely explain the rationales behind the funding decisions made by the two bodies.
@@ 1 Introduction
The STA is interesting because it explicitly acknowledges that OSS is an important part of the world’s critical infrastructures. In other words, the Agency “invests in open digital infrastructure”, which refers to “foundational technologies that enable the creation of other software”, and regarding criticality, people often do not “notice how much our lives depend on digital infrastructure until it stops working(STA, 2024g, ). The OpenSSF’s funding instrument is based on a similar rationale; it seeks to protect “society by catalyzing sustainable security improvements to the most critical open source software projects” (OpenSSF, 2024a, ). Criticality is again one of the main keywords.
@@
To the best of the authors’ knowledge, the paper is the first to examine these two funding bodies and the OSS projects they have funded from a cyber security perspective. In addition to the novelty of the paper’s topic, the paper contributes to two research branches. The first branch is the research on critical infrastructure. As will be argued in the opening Section 2, this branch has often overlooked software’s role in critical infrastructures. In particular, according to a brief but still reasonable literature search, there is very limited—if any—existing work on the relationship between OSS and critical infrastructures.
@@
The second research branch is about the survivability or sustainability of OSS projects (Avelino et al.,, 2019; Chengalur-Smith et al.,, 2010; Decan et al.,, 2020). This branch has focused on a question of what makes open source software projects survive over time. An example of a typical explanatory factor considered in the literature is a project’s ability to attract new developers. However, regarding software quality, including software security with it, human resources are not necessarily what matter the most; software’s growing size, complexity, and technical debt may matter more (Alami et al.,, 2024). Also financial resources may well matter for the security of OSS.
@@
To this and other ends, recent research has also started to examine funding of OSS projects. The paper contributes to this recent OSS funding research and its calls for further research and better engagement (Osborne,, 2024; Sharma,, 2023). In this regard, it should be emphasized that the STA and OpenSSF are by no means the only funding bodies available for OSS projects today. In addition to so-called micro-donations and crowd-sourcing endeavors (Osborne et al.,, 2024), there are new funding initiatives for software companies (FCF,, 2024; OSTIF,, 2024), philanthropic programs (CZI,, 2024), and funding bodies that align with civil liberties (OTF,, 2024). Given the overall funding landscape, including with respect to scientific research (Braun,, 1998; Cocos,, 2020), these and other funding schemes for OSS projects are strategic in their nature; they focus on some particular problem areas. With respect to the STA and OpenSSF, the strategic problem area is cyber security in general and software security in particular. By implication, the paper’s particular focus on the STA and OpenSSF also connects the survivability research branch with the critical infrastructure branch.
@@
Money and cyber security have an intrinsic relationship. In fact, an argument has often been raised that among the foremost reasons for the world’s cyber insecurity is a lack of incentives for companies, societies, and even individuals, including software engineers (Anderson and Moore,, 2006; Halderman,, 2010). This lack of incentives manifests itself in observations that security is not prioritized in the software industry because it is seen as a cost and not a feature (Ryan et al.,, 2023). Although techniques have been proposed to incentivize and prioritize software security (Weir et al.,, 2023), economic factors, such as cost-effectiveness and return-on-investment, have continued to hamper the adoption of new security solutions for software engineering (Ramaj et al.,, 2024). These obstacles are pronounced in the OSS domain already due to the so-called tragedy of the (digital) commons (Greco and Floridi,, 2004; OpenSSF, 2024d, ). Although this classical tragedy is a contested concept (Frischmann et al.,, 2019), it still suits to illustrate a common reasoning; because OSS is openly and publicly available, there is even a littler incentive to financially invest in OSS and its security. This tragedy has been seen as a primary reason behind the new funding initiatives (Sharma,, 2023). The tragedy of the (digital) commons further demonstrates that the fundamental issue is not only about the lack of incentives but actually about misplaced incentives. Because some OSS components are nowadays used in practically all commercial software projects, including those addressing the world’s critical infrastructures, the misplaced incentive to not invest in the security of OSS manifests itself in the world’s overall cyber insecurity, including with respect to digital infrastructures that are vital to the functioning of societies.
@@
The paper’s remainder is structured into five sections. The opening Section 2 outlines the background in more detail with respect to the concept of critical infrastructure and its relation to software. Given this elaboration, the subsequent Section 3 presents and motivates the research questions. Then, Section 4 elaborates the empirical material and the qualitative analysis techniques used to examine the material. Results are presented in Section 5. A concluding discussion is presented in the final Section 6. With these five sections, the paper paves the way for further research on an emerging, relevant new research area, providing also a few insights relevant to practitioners.
2 Background
@@
There is a long-standing debate regarding the concept of infrastructure and what is recognized as critical (Katina and Hester,, 2013; Zebrowski and Sage,, 2017). The jurisprudence of the European Union (EU) provides a good way to briefly illustrate typical definitions and some associated problems. Thus, the EU’s now deprecated Directive 2008/114/EC defined critical infrastructure as “an asset, system or part thereof” that is located in any given member state and “which is essential for the maintenance of vital societal functions, health, safety, security, economic or social well-being of people, and the disruption or destruction of which would have a significant impact”. Although this definition aligns with the classical concept of an asset, which is essential in traditional information security, the societal focus enlarges the scope from mere information to broader assets with vital society-wide functions.
@@
These EU laws have received some criticism in that too many sectors might be covered, expertise might be lacking in some sectors, interoperability and compliance burden may become issues, fragmentation might occur because of sectoral differences and the national adoptions by the member states, and sectors defined as non-critical might receive less resources and effort (Besiekierska,, 2023; Ruohonen, 2024b, ; Vandezande,, 2024; Veigurs et al.,, 2024). A related critical point is that software has received only a limited explicit attention in the laws. Although the EU’s jurisprudence recognizes information technology as a distinct critical infrastructure of its own, it is important to emphasize that also all of the other critical infrastructures noted are dependent on software. Neither banking nor a central government can nowadays function without software. Nor are there space technologies without software. In fact, it can be argued that a new “software-defined” paradigm is emerging—or has already emerged, and this paradigm also amplifies security risks due to software’s insecurity and other problems (Wu,, 2024). This point justifies also the STA’s and OpenSSF’s overall rationales.
@@ 3 Research Questions
The discussion in the previous section is enough to motivate the first research question (RQ):
@@
    RQ.1: What kind of OSS the STA and OpenSSF have funded and thus considered as critical?
An answer to the question sheds important light on what kind of OSS is seen as critical by a governmental agency and a not-for-profit OSS foundation in terms of cyber and software security. Indirectly, the question also aligns with existing work on how to identify critical infrastructures (Di Mauro et al.,, 2010; Katina and Hester,, 2013). In this regard, it can be also remarked that the OpenSSF, 2024d has a specific working group trying to solve the issue of how to identify critical open source software projects. For the present purposes, an answer to RQ.1 can be reflected also from a broader perspective; if some particular type of OSS stands out, it may be that the EU’s existing legal frameworks have overlooked some important aspects when operating with the criticality concept.
In this regard, a further EU law can be noted: Regulation (EU) 2024/2847, which is commonly known as the Cyber Resilience Act (CRA). It lays down some cyber security requirements for commercially supported OSS projects, including particularly with respect to vulnerability disclosure and associated coordination. Open source foundations, which often fund open source projects, are also mandated to develop policies for handling vulnerabilities and fostering secure software development. In addition, commercial vendors who integrate OSS into their products are mandated to notify the associated OSS projects about vulnerabilities they have discovered or been made aware of. While keeping in mind potential unintended consequences (Ruohonen, 2024b, ; Ryan et al.,, 2023), these mandates can be seen to implicitly address the noted interlinking aspect of critical infrastructures. The EU has also established new funding schemes to support the implementation and adoption of the new cyber security laws in the public and private sectors (Ruohonen, 2024b, ; Veigurs et al.,, 2024). However, these funding instruments exclude OSS projects; hence, an answer to RQ.1 can also help to understand the point about regulators and political decision-makers perhaps overlooking important aspects related to critical infrastructures and their development, maintenance, and operation.
Regarding grants and funding in general, the following research question is always worth asking:
@@
The question can be motivated by the extensive research on choosing OSS components for a software project, whether commercial or open source. This research area is closely related to the survivability and sustainability aspects noted in the introduction. The list of factors to consider is extensive already in terms of security alone. The examples include reported vulnerability counts for a component (Cox,, 2019), the centrality of a component in a dependency network and the trust placed upon its maintainers (Zimmermann et al.,, 2019), warnings from static analyzers (Ruohonen et al.,, 2024) and other software testing tools, and even recommender systems using these and other factors for recommendations (Nguyen et al.,, 2018; Ruohonen,, 2019). By and large, however, the existing research has operated with quantified metrics that can be easily extracted from OSS projects. To this end, an answer to RQ.3 addresses a notable limitation in existing research because it may well be that a rationale for funding has only a little to do with the conventional, mechanical metrics. This point entwines with previous observations according to which developers mostly concentrate on functionality and not security when choosing OSS components and dependencies (Pashchenko et al.,, 2020). Here too the incentives noted in the introduction are visible. In any case, an answer to RQ.3 can also be beneficial to OSS projects and their developers in formulating competitive funding proposals. In this regard, RQ.3 further aligns with existing research on guidelines for writing grant applications in the domain of software projects (Strasser et al.,, 2022). In addition to the types of software and the rationales for funding decisions, it is relevant to know what kind of cyber security the funded projects have addressed:
@@ 4 Materials and Methods
The data is based on all funding cases that were available from the OpenSSF, 2024a’s () and STA, 2024g’s () websites during the start of the data collection on 27 November, 2024. The data includes all funding cases listed on the websites, including those without longer funding descriptions. In total, $n=24$ and $n=44$ funding cases are covered from the OpenSSF and STA, respectively.
The paper’s qualitative methodology is based on thematic analysis. Somewhat like topic modeling in quantitative research, this classical qualitative method seeks to find latent themes present in a corpus. An inductive logic was used for the identification of themes and their refinement. By following guidelines (Vaismoradi et al.,, 2013), the thematic analysis was carried out in five steps. The first step was about familiarizing with the qualitative data, meaning that the funding cases were read and reread, and initial notes were taken. Based on this familiarizing, the second step involved a generation of initial codes for the potential themes. Then, the third and fourth steps were about searching and identifying themes based on the codes generated, and then collating overlapping themes due to a goal of parsimonious answers to the five research questions specified. The fifth and final step involved defining and naming the final themes. Although the result are summarized with concise, code-like themes together with their occurring frequencies, the common pitfalls of mere categorization and plain quantification (Ruohonen, 2024a, ; Saunders et al.,, 2023) are addressed by illustrating the themes with further elaborations and contemplations relevant to further research.
@@
It can be noted that thematic analysis is often triangulated with other qualitative methodologies, including particularly what is known as grounded theory (Ruohonen, 2024a, ). However, a step toward deductive reasoning with prior theoretical premises was omitted on the grounds that the existing OSS funding research is too scant for developing sound theoretical building blocks. Though, regarding grounded theory, which essentially assumes that a theory can be found from data—it is grounded there, the final Section 6 discusses also theoretical implications of the qualitative results reached. Grounded theory provides also means to further elaborate the coding process behind the thematic analysis: the initial identification of themes involved open coding, which resulted many themes, which were then refined and reduced by the means of axial coding, and selective coding was finally used for collating and sharpening the themes from the axial coding (Stol et al.,, 2016; Williams and Moser,, 2019). In addition to this procedural coding logic, it is important to emphasize that the paper’s all authors were involved in the analysis.
@@
Thus, to further elaborate the collaborative side of the thematic analysis, the initial analysis was carried out by the paper’s corresponding author. Then, the paper’s other collaborating authors independently reviewed, case by case, the results from the initial analysis. They were given a chance to disagree and propose corrections, including the introduction of new themes. With respect to disagreements and diverging themes, the corresponding author made the final decision about a theme by evaluating the reviews of the collaborating authors. Although inter-rater reliability measures were omitted for brevity, this collaborative process follows the recent suggestion to work sequentially in qualitative coding (Díaz et al.,, 2023). Even though independent and blinded coding was not done, the collaborative approach can be argued to still improve the rigor and trustworthiness of the qualitative results.

### v2

@@ Abstract
Context: Many open source software (OSS) projects need more human resources for maintenance, improvements, and sometimes even their survival. This need allegedly applies even to vital OSS projects that can be seen as being a part of the world’s critical infrastructures. To address this resourcing problem, new funding instruments for OSS projects have been established in recent years. Objectives: The paper examines two such funding bodies for OSS and the projects they have funded. The focus of both funding bodies is on software security and cyber security in general. Methods: The methodology is based on qualitative thematic analysis. Results: Particularly OSS supply chains, network and cryptography libraries, programming languages, and operating systems and their low-level components have been funded and thus seen as critical in terms of cyber security by the two funding bodies. Conclusions: In addition to the qualitative results presented, the paper makes a contribution by connecting the research branches of critical infrastructure and sustainability of OSS projects. A further contribution is made by connecting the topic examined to recent cyber security regulations. Finally, an important argument is raised that neither cyber security nor sustainability alone can entirely explain the rationales behind the funding decisions made by the two bodies.
@@ 1 Introduction
For the paper’s purposes, the STA is interesting because it explicitly acknowledges that OSS is an important part of the world’s critical infrastructures. In other words, the Agency “invests in open digital infrastructure”, which refers to “foundational technologies that enable the creation of other software”, and regarding criticality, people often do not “notice how much our lives depend on digital infrastructure until it stops working(STA, 2024i, ). The OpenSSF’s funding instrument is based on a similar rationale; it seeks to protect “society by catalyzing sustainable security improvements to the most critical open source software projects” (OpenSSF, 2024a, ). Criticality is again one of the main keywords.
@@
To the best of the authors’ knowledge, the paper is the first to examine these two funding bodies and the OSS projects they have funded from a cyber security perspective. In addition to the novelty of the paper’s topic, the paper contributes to two research branches. The first branch is the research on critical infrastructure. As will be argued in the opening Section 2, this branch has often overlooked software’s role in critical infrastructures. In particular, there appears to be very limited existing work on the relationship between OSS and critical infrastructures.
@@
The second research branch is about the survivability or sustainability of OSS projects (Alami et al.,, 2024; Avelino et al.,, 2019; Chengalur-Smith et al.,, 2010; Decan et al.,, 2020; Tulili et al.,, 2025). This branch has focused on a question of what makes open source software projects survive over time. An example of a typical explanatory factor considered in the literature is a project’s ability to attract new developers. However, regarding software quality, including software security with it, human resources are not necessarily what matter the most; software’s growing size, complexity, and technical debt may matter more (Alami et al.,, 2024). Also financial resources may well matter for the security of open source software.
@@
To this and other ends, recent research has also started to examine funding of OSS projects. The paper contributes to this recent OSS funding research and its calls for further research and better engagement (Osborne,, 2024; Kula and Treude,, 2025; Sharma,, 2023; Tsakpinis and Pretschner,, 2025). In this regard, it should be emphasized that the STA and OpenSSF are by no means the only funding bodies available for OSS projects today. In addition to so-called micro-donations and crowd-sourcing endeavors (Osborne et al.,, 2024; Tsakpinis and Pretschner,, 2025), there are new funding initiatives for companies to support open source software projects (FCF,, 2025; OSTIF,, 2025), philanthropic programs (CZI,, 2025; NLnet Foundation,, 2025), and funding bodies that align with civil liberties (OTF,, 2025). Given the overall funding landscape, including with respect to scientific research (Braun,, 1998; Cocos,, 2020), these and other funding schemes for OSS projects are strategic in their nature; they focus on some particular problem areas. With respect to the STA and OpenSSF, the strategic problem area is cyber security in general and software security in particular. By implication, the paper’s particular focus on the STA and OpenSSF also connects the survivability research branch with the critical infrastructure branch.
@@
Money and cyber security have an intrinsic relationship. In fact, an argument has often been raised that among the foremost reasons for the world’s cyber insecurity is a lack of incentives for companies, societies, and even individuals, including software engineers (Anderson and Moore,, 2006; Halderman,, 2010). This lack of incentives manifests itself in observations that security is not prioritized in the software industry because it is seen as a cost and not a feature (Ryan et al.,, 2023). Although techniques have been proposed to incentivize and prioritize software security (Weir et al.,, 2023), economic factors, such as cost-effectiveness and return-on-investment, have continued to hamper the adoption of new security solutions for software engineering (Ramaj et al.,, 2024). These obstacles are pronounced in the OSS domain already due to the so-called tragedy of the (digital) commons (Greco and Floridi,, 2004; OpenSSF, 2024e, ). Although this classical tragedy is a contested concept (Frischmann et al.,, 2019), it serves to illustrate a common reasoning; because OSS is openly and publicly available, there is even a less incentive to financially invest in OSS and its security. This tragedy has been seen as a primary reason behind the new funding initiatives (Sharma,, 2023). The tragedy of the (digital) commons further demonstrates that the fundamental issue is not only about the lack of incentives but actually about misplaced incentives. Because some OSS components are nowadays used in practically all commercial software projects, including those addressing the world’s critical infrastructures, the misplaced incentive to not invest in the security of OSS manifests itself in the world’s overall cyber insecurity, including with respect to digital infrastructures that are vital to the functioning of societies.
@@
The paper’s remainder is structured into five sections. The opening Section 2 motivates the background in more detail with respect to the concept of critical infrastructure and its relation to software. Given this elaboration, the subsequent Section 3 presents and motivates the research questions. Then, Section 4 elaborates the empirical material and the qualitative analysis techniques used to examine the material. Results are presented in Section 5. A concluding discussion is presented in the final Section 6. With these remaining five sections, the paper paves the way for further research on an emerging, relevant new research area, providing also a few insights relevant to practitioners, including OSS projects seeking funding, the two funding bodies’ staff, and policy-makers more generally.
2 Motivation
@@
There is a long-standing debate regarding the concept of infrastructure and what is recognized as critical (Eckhardt and Kotovskaia,, 2023; Katina and Hester,, 2013; Zebrowski and Sage,, 2017). The jurisprudence of the European Union (EU) provides a good way to briefly illustrate typical definitions and some associated problems. Thus, the EU’s now deprecated Directive 2008/114/EC defined critical infrastructure as “an asset, system or part thereof” that is located in any given member state and “which is essential for the maintenance of vital societal functions, health, safety, security, economic or social well-being of people, and the disruption or destruction of which would have a significant impact”. Although this definition aligns with the classical concept of an asset, which is essential in traditional information security, the societal focus enlarges the scope from mere information to broader assets with vital society-wide functions.
@@
These EU laws have received some criticism in that too many sectors might be covered, expertise might be lacking in some sectors, interoperability and compliance burden may become issues, fragmentation might occur because of sectoral differences and the national adoptions by the member states, and sectors defined as non-critical might receive less resources and effort (Besiekierska,, 2023; Ruohonen,, 2024; Vandezande,, 2024; Veigurs et al.,, 2024). A related critical point is that software has received only a limited explicit attention in the laws. Although the EU’s jurisprudence recognizes information technology as a distinct critical infrastructure of its own, it is important to emphasize that also all of the other critical infrastructures noted are dependent on software. Neither banking nor a central government can nowadays function without software. Nor are there space technologies without software. In fact, it can be argued that a new “software-defined” paradigm is emerging—or has already emerged, and this paradigm also amplifies security risks due to software’s insecurity and other problems (Wu,, 2024). This point justifies also the STA’s and OpenSSF’s overall rationales.
@@ 3 Research Questions
The discussion in the previous section motivates the first research question (RQ):
@@
    RQ.1: What kind of OSS have the STA and OpenSSF funded and thus considered as critical?
An answer to RQ.1 sheds important light on what kind of OSS is seen as critical by a government-backed agency and a not-for-profit OSS foundation in terms of cyber and software security. Indirectly, the question also aligns with existing work on how to identify critical infrastructures (Di Mauro et al.,, 2010; Katina and Hester,, 2013). In this regard, it can be also remarked that the OpenSSF, 2024e has a specific working group trying to solve the issue of how to identify critical open source software projects. For the present purposes, an answer to RQ.1 can be reflected also from a broader perspective; if some particular type of OSS stands out, it may be that the EU’s existing legal frameworks have overlooked some important aspects when operating with the criticality concept.
In this regard, a further EU law can be noted: Regulation (EU) 2024/2847, which is commonly known as the Cyber Resilience Act (CRA). Unlike in the NIS2 directive, criticality in this regulation is not approached through sectors but through products (Eckhardt and Kotovskaia,, 2023). It also specifies some new cyber security requirements for commercially supported OSS projects, including with respect to vulnerability disclosure and associated coordination (Ruohonen and Timmers,, 2024). Open source foundations, which often fund open source projects, are also mandated to develop policies for handling vulnerabilities and fostering secure software development. In addition, commercial vendors who integrate OSS into their products are mandated to notify the associated OSS projects about vulnerabilities they have discovered or been made aware of. Meeting the CRA’s essential cyber security requirements, which have been recently analyzed (Ruohonen et al.,, 2025), in the commercial software industry necessitates dealing also with OSS because many—if not most—commercial software projects use at least some OSS components. For instance, among the essential requirements is an obligation to ship products without known vulnerabilities. This legal requirement applies also to all OSS components embedded to a product.
While keeping in mind potential unintended consequences (Ruohonen,, 2024; Ryan et al.,, 2023), these legal mandates can be seen to implicitly address the noted interlinking aspect of critical infrastructures. The EU has also established new funding schemes to support the implementation and adoption of the new cyber security laws in the public and private sectors (Ruohonen and Timmers,, 2025; Veigurs et al.,, 2024). However, these funding instruments seem to exclude OSS projects; hence, an answer to RQ.1 can also help to understand the point about regulators and political decision-makers perhaps having overlooked important aspects related to critical infrastructures and their development, maintenance, and operation.
Regarding grants and funding in general, the following research question is worth asking:
@@
The question can be motivated by the extensive research on choosing OSS components for a software project, whether commercial or open source. This research area is closely related to the survivability and sustainability aspects noted in the introduction. The list of factors to consider is extensive already in terms of security alone. The examples include reported vulnerability counts for a component (Cox,, 2019), the centrality of a component in a dependency network and the trust placed upon its maintainers (Zimmermann et al.,, 2019), and even recommender systems using these and other factors for recommendations (Nguyen et al.,, 2018). By and large, however, the existing research has operated with quantified metrics that can be easily extracted from OSS projects. To this end, an answer to RQ.3 addresses a notable limitation in existing research because it may well be that a rationale for funding has only a little to do with the conventional, mechanical metrics. This point entwines with previous observations according to which developers mostly concentrate on functionality and not security when choosing OSS components and dependencies (Pashchenko et al.,, 2020). Here too the incentives noted in the introduction are visible. In any case, an answer to RQ.3 can also be beneficial to OSS projects and their developers in formulating competitive funding proposals. In this regard, RQ.3 further aligns with existing research on guidelines for writing grant applications in the domain of software projects (Strasser et al.,, 2022). In addition to the types of software and the rationales for funding decisions, it is relevant to know what kind of cyber security the funded projects have addressed:
@@ 4 Materials and Methods
The data is based on all funding cases that were available from the OpenSSF, 2024a’s () and STA, 2024i’s () websites during the start of the data collection on 27 November, 2024. The data includes all funding cases listed on the websites, including those without longer funding descriptions. In total, $n=24$ and $n=44$ funding cases are covered from the OpenSSF and STA, respectively.
The paper’s qualitative methodology is based on thematic analysis. Somewhat like topic modeling in quantitative research, this classical qualitative method seeks to find latent themes present in a corpus. An inductive logic was used for the identification of themes and their refinement. By following guidelines (Terry et al.,, 2017; Vaismoradi et al.,, 2013), the thematic analysis was carried out in five steps. The first step was about familiarizing with the qualitative data, meaning that the funding cases were read and reread, and initial notes were taken. Based on this familiarizing, the second step involved a generation of initial codes for the potential themes. Then, the third and fourth steps were about searching and identifying themes based on the codes generated, and then collating overlapping themes due to a goal of parsimonious answers to the five research questions specified. The fifth and final step involved defining and naming the final themes. Although the result are summarized with concise, code-like themes together with their occurring frequencies, the common pitfalls of mere categorization and plain quantification (Saunders et al.,, 2023) are addressed by illustrating the themes with further elaborations and contemplations relevant to further research.
@@
It can be noted that thematic analysis is sometimes triangulated with other qualitative methodologies, including what is known as grounded theory. However, a step toward deductive reasoning with prior theoretical premises was omitted on the grounds that the existing OSS funding research is too scant for developing sound theoretical building blocks. Though, regarding grounded theory, which essentially assumes that a theory can be found from data—it is grounded there, the final Section 6 discusses also theoretical implications of the qualitative results reached. Grounded theory provides also means to further elaborate the coding process behind the thematic analysis: the initial identification of themes involved open coding, which resulted in many themes, which were then refined and reduced by the means of axial coding, and selective coding was finally used for collating and sharpening the themes from the axial coding (Stol et al.,, 2016; Williams and Moser,, 2019). What separates the coding done from grounded theory is that no explicit theorization was done alongside the analysis and the themes were not taken as latent (cf. Terry et al.,, 2017). In addition to the procedural coding logic, it is important to emphasize that all the paper’s authors were involved in the analysis.
It is also worth emphasizing that the coding was rather straightforward due to the relatively short descriptions provided for the funding decisions. In general, the STA tends to provide a little more information than the OpenSSF, but even that information typically only amounts to a few short paragraphs. Both funding bodies describe their decisions in unstructured formats. In addition to the slight variance between the funding bodies, there is a variance between individual funding decisions; some are described shortly, while others are elaborated in detail. This variance justifies also the qualitative approach because a quantitative approach might be biased due to the variance. The simplicity of the final, collated themes eased the coding and its robustness. For instance, it is rather trivial to determine whether a funded project is about operating systems (e.g., OpenSSF, 2024b, ), logging libraries (STA, 2024d, ), desktop environments (STA, 2024c, ), or programming language ecosystems (e.g., OpenSSF, 2024c, ; STA, 2024b, ). As can be deduced from Fig. 1, most of the funding decisions and the themes for these do not require complicated reasoning.
Figure 1: An Example of a Funding Decision (STA, 2024d, )
@@
Thus, to further elaborate the collaborative side of the thematic analysis, the initial analysis was carried out by the paper’s first author. Then, the paper’s other collaborating authors independently reviewed, case by case, the results from the initial analysis. They were given a chance to disagree and propose corrections, including the introduction of new themes. With respect to disagreements and diverging themes, the first author made the final decision about a theme by evaluating the reviews of the collaborating authors. Although inter-rater reliability measures were omitted, this collaborative process follows the recent suggestion to work sequentially in qualitative coding (Díaz et al.,, 2023). The collaborative approach is also known as researcher triangulation in qualitative research (Farquhar et al.,, 2020). Regarding inter-rater reliability measures, the omission of these can be justified on the grounds that only the final themes are relevant and thus reported, the prior knowledge of the authors varied regarding the subject matter, and the funding decisions were rather straightforward to interpret and categorize (cf. McDonald et al., 2019; and more generally Terry et al., 2017 who argue against the measures altogether in qualitative research). Even though independent and blinded coding was not done, the collaborative approach can be argued to still improve the rigor and trustworthiness of the qualitative results. Though, with the earlier Fig. 1 in mind, only a few decisions prompted some initial disagreements.